OpenStudy (anonymous):

Find the surface area of a cone with radius of 8.8 cm and slant height of 7.1 cm to the nearest square unit. Use 3.14 for pi.

OpenStudy (anonymous):

the equation you will use here is: SA = πr^2 + πrl

OpenStudy (anonymous):

SA = πr^2 + πrl SA = π(8.8)^2 + π(8.8)(7.1) SA = 77.44π + 62.48π SA = 139.92π SA ≈ 439.57

OpenStudy (anonymous):

Got it! Thanks a lot both of you!